Problem 1

Calculate and interpret the correlation coefficient of the two variables below.

Person

Hand

Height


A

17

150


B

15

154


C

19

169


D

17

172


E

21

175



1
Expert's answer
2021-06-24T18:32:44-0400

Pearson Correlation coefficient



"r = \frac{\sum(x_i - \bar x) (y_i - \bar y)}{\sum(x_i - \bar x)^2 \sum(y_i - \bar y)^2}"

Let "x\ be\ hand \ and \ y\ be\ height"


"\bar x =\frac{17+15+19+17+21}{5}= 17.8"

"\bar y = \frac{150+154+169+172+175}{5}=164"


"\sum(x_i - \bar x) (y_i - \bar y)=(17-17.8)(150-164)+(25-17.8) (154-164)+(19-17.8)(169-164)+(17-17.8)(172-164)+(21-17.8)(175-164)""=-26"


"\sum(x_i - \bar x)^2=(17-17.8)^2+(25-17.8)^2+(19-17.8)^2+(17-17.8)^2+(21-17.8)^2""=64.8"


"\sum (y_i - \bar y)^2=(150-164)^2+ (154-164)^2+(169-164)^2+(172-164)^2+(175-164)^2""=506"

Therefore


"r=\frac{-26}{506*64.8}=-0.000793"

Since we have 10 data points, the degrees of freedom are


"10-2=8"

The critical value for the correlation coefficient with 8 degrees of freedom is

"\underset{-}{+} 0.632." We conclude that the correlation is significant if it is greater than the critical value. In our case, "-0.000793 > - 0.632" therefore we can conclude that there is a significant correlation between hand and height